OpenStudy (vane11):

Can someone explain to me how to calculate r^2 in statistics and how it's useful? My notes aren't really helping me understand and I have a test today :(

OpenStudy (amistre64):

sounds like a correlation coefficient ....

OpenStudy (vane11):

well I have the definition of it, so I understand it's purpose or what it does, but that's it

OpenStudy (amistre64):

yeah, something about an explained variation

OpenStudy (amistre64):

the formula looks a bit convoluted, but seems to be a pattern repeated 3 times

OpenStudy (amistre64):

\[r=\frac{n~xy-(x)(y)}{\sqrt{n~x^2-(x)^2}~\sqrt{n~y^2-(y)^2}}\]

OpenStudy (amistre64):

a simplified notation addresses something called a zscore for a sample value .... but i got no idea what that would be talking about

OpenStudy (vane11):

well I have r^2=(y-y-hat)2 ....

OpenStudy (vane11):

i was given like 2, 3 equations so idk if there's one that's better than the others or if they're all equally correct

OpenStudy (amistre64):

i havent delved into the chapters in my book that cover it, but it looks like excel would be very useful for this setup

OpenStudy (amistre64):

or a table to organize the stuff x y xy x^2 y^2 .. .. .. .. .. sums||

OpenStudy (amistre64):

and maybe a spot to remember the n :)

OpenStudy (amistre64):

spose for brevity that x={2,4,5} and y = {1,3,3} x y xy x^2 y^2 2 1 2 4 1 4 3 12 16 9 5 3 15 25 9 sums|| 11 10 29 45 19 and there are 3 rows so n=3 \[r=\frac{3(29)-(5)(3)}{\sqrt{3(45)-11}\sqrt{3(9)-3}}\] \[r=\frac{72}{\sqrt{124}\sqrt{24}}\] \[r^2=\frac{72^2}{124*24}\]

OpenStudy (amistre64):

thats spose to be (11)(10) not (5)(3) up top

OpenStudy (amistre64):

and 3(19) on the right under there not 3(9) .... lol

OpenStudy (amistre64):

other than that ..

OpenStudy (amistre64):

\[r=\frac{n~\sum(xy)-\sum x \sum y}{\sqrt{n~\sum(x^2)-(\sum x)^2}~\sqrt{n~\sum(y^2)-(\sum y)^2}}\] \[r=\frac{3(29)-(11)(10)}{\sqrt{3(45)-11}\sqrt{3(19)-10}}\] oy thats was painful :)